OpenStudy (anonymous):

A street light is on top of a 10 foot pole. A person who is 4 feet tall walks away from the pole at a rate of 6 feet per second. At what speed is the length of the person’s shadow growing?

OpenStudy (anonymous):

did you learn APs and GPs yet?

OpenStudy (anonymous):

no :( I don't know what those are

OpenStudy (anonymous):

(y + x)/10 = y/6 6(y + x) = 10y 6y + 6x = 10y 6x = 4y 6(dx/dt) = 4(dy/dt) (6/4)(3) = dy/dt 18/4 = dy/dt 9/2 ft/s = dy/dt